OpenStudy (anonymous):

A particular model of walkie-talkie can broadcast in a circular area. The radius of the broadcast area is 10,000 feet. Find the area of this circle to the nearest square foot. Use 3.14 for p.

OpenStudy (shane_b):

Ops...make that: \[A_{circle}=\pi r^2\] I just did a cylinder area a moment ago and pasted the wrong thing :)

OpenStudy (anonymous):

still confused

OpenStudy (shane_b):

\[\pi=3.15\]\[r=10,000ft\]\[A=\pi r^2\]Plug in the value for r and pi and you'll get your answer.

OpenStudy (anonymous):

so 10,000 and 3.14 and what do i do with it?

OpenStudy (shane_b):

\[r^2=10000^2\]Now just multiply that times 3.14 to get the area.

OpenStudy (anonymous):

314,000,000 ft2

OpenStudy (shane_b):

Sorry...yep, that's correct.